# Modality Collapse as Mismatched Decoding: Information-Theoretic Limits of Multimodal LLMs

### v1

@@ A.2 GMI-Wasserstein bound
We now bound the GMI degradation when the decoder sees modal representations $P_{\!M}$ instead of the text representations $P_{\!T}$ it was trained on. The intuition is simple: if modal representations are close to text representations, the decoder should perform nearly as well on them. Two properties of the decoder determine how “close” is close enough: how sensitive it is to small shifts in its input, and how large the space of possible representations is.
@@
There exists $L_{\log}>0$ such that for all $(c,y)$ and all $z,z^{\prime}$ in the typical-state region $\mathcal{Z}$: $|\ell_{\psi}(c,z,y)-\ell_{\psi}(c,z^{\prime},y)|\leq L_{\log}\|z-z^{\prime}\|$.
That is, if you nudge a representation by a small amount, the decoder’s log-probability changes by at most $L_{\log}$ times that amount. The constant $L_{\log}$ captures how sensitive the decoder is to shifts in its input. Neural networks with smooth activations satisfy this locally; we estimate $L_{\log}$ empirically via gradient norms (§5.5).
@@
This says the representations that actually occur in practice live in a bounded region, not scattered arbitrarily through $\mathbb{R}^{d}$. LayerNorm ensures this by projecting all hidden states onto a hypersphere of radius $\sqrt{d}$.
@@
$|\mathrm{GMI}_{P_{\!M}}(q_{\psi})-\mathrm{GMI}_{P_{\!T}}(q_{\psi})|\leq\big(1+e^{L_{\log}\cdot D}\big)\cdot L_{\log}\cdot\mathbb{E}_{(C,Y)}\!\left[W_{\!1}\!\big(P_{\!M}(Z|C,Y),\,P_{\!T}(Z|C,Y)\big)\right].$ (4)
Reading the bound. In words:
@@
The gap between what the decoder could extract from text representations and what it can extract from modal representations is controlled by how far apart those distributions are, amplified by how sensitive the decoder is. Crucially, $W_{\!1}$ measures distributional distance, not just the distance between means: aligning the average modal representation with the average text representation is not enough. If the modal distribution has different shape, spread, or tails, the exponential penalty factor $(1+e^{L_{\log}D})$ catches them. This is why text-aligned encoders reduce the bound (they project onto text-like directions, reducing both the distance and the shape mismatch) while simple mean-centering would not.
The prefactor $(1+e^{L_{\log}D})$ arises from the “competition term” in the proof (see below). In the ambient space this prefactor is vacuously large ($L_{\log}\cdot D$ ranges from 10 to 150); the support-restricted bound (§A.3) tightens it using the effective support diameter.
@@ A.2.1 Proof of Theorem 2
Proof overview. The GMI has two pieces: (A) how well the decoder scores the correct representation, and (B) how well it scores random alternatives. We need to show that switching from text to modal representations does not change either piece by too much. For piece (A), the argument is simple: if the decoder’s scores change smoothly (Lipschitz assumption), then moving representations a small distance changes expected scores by a small amount. For piece (B), the argument is harder because piece (B) involves an exponential of the scores, which amplifies small changes. The bounded-diameter assumption (representations live in a finite region) keeps this amplification under control.
Setup. Write $\ell(z)\equiv\ell_{\psi}(c,z,y)$ for fixed $(c,y)$. The GMI functional decomposes as $\Gamma(P)=A(P)-B(P)$ where:
@@
We bound $|A(P_{\!M})-A(P_{\!T})|$ and $|B(P_{\!M})-B(P_{\!T})|$ separately.
Step 1: Direct term (how much does the correct representation’s score change?). We want to bound how much the expected score changes when we switch from $P_{\!T}$ to $P_{\!M}$. The Kantorovich–Rubinstein theorem (Kantorovich & Rubinstein, 1958; Villani, 2009) says: if a function changes by at most $L$ when you move its input by 1 unit ($L$-Lipschitz), then the difference in its expected value under two distributions is at most $L$ times the Wasserstein distance between those distributions. Since $\ell$ is $L_{\log}$-Lipschitz (Assumption 2):
@@
In words: if modal and text representations are close (small $W_{\!1}$) and the decoder’s scores change smoothly (small $L_{\log}$), then the expected score barely changes.
Step 2: Competition term (how much does the random-alternative score change?). This is harder because $B(P)=\log\mathbb{E}[e^{\ell(Z)}]$ involves an exponential of the score, which amplifies small differences. A coupling $\pi$ is a way of pairing up samples from $P_{\!M}$ and $P_{\!T}$ (imagine lining up modal representations next to the text representations they most resemble). For each pair $(Z_{M},Z_{T})$, define $\Delta\ell=\ell(Z_{M})-\ell(Z_{T})$: the difference in score between the paired representations. By the Lipschitz assumption, $|\Delta\ell|\leq L_{\log}\cdot\|Z_{M}-Z_{T}\|$, and by the bounded diameter (Assumption 3), $|\Delta\ell|\leq L_{\log}\cdot D$.
@@
The first factor $e^{\ell(Z_{T})}/m(P_{\!T})$ is a probability weight under $P_{\!T}$ (it sums to 1). The second factor $e^{\Delta\ell}$ captures how the shift from $Z_{T}$ to $Z_{M}$ changes the score. Since $|e^{x}-1|\leq|x|\cdot e^{|x|}$ for bounded $|x|\leq L_{\log}D$:
@@
Using $|\log(1+x)|\leq|x|$ for $|x|<1$ (applicable when $W_{\!1}$ is small enough that the ratio stays near 1):
@@
The $e^{L_{\log}D}$ prefactor is the price of the exponential: the competition term is more sensitive to distributional shift than the direct term because it involves $e^{\ell}$ rather than $\ell$. This is analogous to how compound interest amplifies small rate changes: a 1% shift in the base rate produces a much larger change in the compounded amount when the compounding period is long.
Step 3: Combine. The total GMI change is bounded by the sum of the direct and competition terms:
@@
The competition term dominates: it is $e^{L_{\log}D}$ times larger than the direct term. This is why the bound’s prefactor is large, and why the support-restricted corollary (which reduces $D$) is practically important.
Step 4: Average over $(C,Y)$. Everything above was for a fixed prompt $C$ and target $Y$. By Assumption 1 ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), the same prompts and targets appear under both laws, so we simply average over all $(C,Y)$ pairs to obtain the final bound.
@@ A.3 Support-restricted bound
Support-restricted bound. If $P_{\!M}(Z|c,y)$ and $P_{\!T}(Z|c,y)$ are both supported within a region $\mathcal{S}\subseteq\mathcal{Z}$ with $D_{\mathcal{S}}=\mathrm{diam}(\mathcal{S})\leq D$, then $D$ may be replaced by $D_{\mathcal{S}}$ in Theorem 2.
Proof. If both marginals are supported in $\mathcal{S}$, any coupling $\pi$ has support on $\mathcal{S}\times\mathcal{S}$, so $\|Z_{M}-Z_{T}\|\leq D_{\mathcal{S}}$ for $\pi$-a.e. $(Z_{M},Z_{T})$. Step 2 above then bounds $|\Delta\ell|\leq L_{\log}\cdot D_{\mathcal{S}}$, giving the tighter prefactor.
Why this matters. The ambient diameter $D$ makes the prefactor $(1+e^{L_{\log}D})$ astronomically large ($L_{\log}\cdot D$ ranges from 10 to 150). But modal and text representations often occupy a much smaller effective region. We estimate the effective support diameter $D_{\text{eff}}$ via the participation ratio of the pooled representation covariance (§5.5), yielding $L_{\log}\cdot D_{\text{eff}}\in[3,7]$ and a moderate prefactor ($\sim$20–1100).
A.4 Probe-decoder asymmetry
@@
Proof. Identical to Step 1 of the Theorem 2 proof, replacing $\ell_{\psi}$ with $\log h$ and $L_{\log}$ with $L_{h}$. Crucially, there is no competition term for probes (no random-coding argument), so the exponential prefactor does not appear.
The key asymmetry. A linear probe is a simple instrument: one matrix multiplication followed by a softmax. Its sensitivity to distributional shift ($L_{h}$) is just the size of that matrix’s weights. The LLM decoder is a deep transformer with billions of parameters, and its sensitivity ($L_{\log}$) reflects that depth. Empirically, $L_{\log}/L_{h}\approx 30\times$5.5).
Now consider the same distributional shift (modal representations replacing text representations). The probe’s bound says degradation $\leq L_{h}\cdot W_{\!1}$ (small). The decoder’s bound says degradation $\leq(1+e^{L_{\log}D})\cdot L_{\log}\cdot W_{\!1}$ (large). Same shift, dramatically different impact. This is the formal basis for the “present but inaccessible” phenomenon: a probe can still read the information because it is a simple, low-sensitivity instrument, but the decoder’s complex scoring rule is far more sensitive to the distributional mismatch and cannot use it.
@@ B.2 Lipschitz estimation
$L_{\log}$ estimated via the 2-nearest-neighbor gradient norm method over n=1,000n=1{,}000 samples. For each sample $z_{i}$, we compute $\|\nabla_{z}\log q_{\psi}(y_{i}|z_{i},c_{i})\|$ via autograd. The 95th percentile of the resulting distribution serves as the $L_{\log}$ estimate.
@@ B.3 Extraction prompts
During representation extraction, each model receives a minimal, task-neutral prompt to avoid biasing hidden states toward any particular downstream task. Representations are extracted from intermediate hook points, not from the generated output.
@@ B.4 LoRA experiment details
LoRA configuration. Rank r=16r{=}16, $\alpha{=}32$, dropout $0.05$, target modules: q_proj, k_proj, v_proj, o_proj (same configuration as the ALME LoRA used as the negative example).
@@
Loss is computed only on the assistant response tokens; all input/prompt tokens are masked with label 100-100.
@@
Baseline evaluation. Base Ultravox (no LoRA) achieves $17.3\%$ generation accuracy on the held-out split (near chance for 6 classes), defaulting to “neutral” for most inputs.
@@ B.5 Full probe accuracy tables
Table 5 reports the full probe accuracies (mean $\pm$ std over 5 seeds) at all four hook points for every model–dataset–information-type combination. The key pattern to observe: probe accuracy for non-text information types remains well above chance even at the LLM’s final layer, confirming that the information survives the decoder’s processing. It is the decoder’s use of this information, not its presence, that fails.
@@ B.6 Information retention overview
Table 6 summarizes the information retention pattern across all five models. Retention is computed as the ratio of probe accuracy at the LLM’s final layer to probe accuracy at the adapter output, normalized by chance. Values above 100% indicate the LLM amplifies the signal; values below 100% indicate degradation. The key pattern: non-aligned encoders show asymmetric behavior (lexical content is amplified while speaker identity collapses), whereas text-aligned encoders retain or slightly improve all information types, but only because the encoder has already discarded non-textual information.
@@ B.9 Audio-token-only pooling
Table 9 shows the result: audio-only pooling produces nearly identical probe accuracies across all datasets and information types, with all deltas within $\pm$2.1% and most under 1%. Restricting to audio tokens does not help, and in several cases slightly hurts. This is because the LLM’s self-attention layers redistribute information across all sequence positions as they process the input. By layer 16, audio-derived information has spread to text prompt tokens as well. Pooling over the full sequence therefore captures more of the available signal, not less. This validates our use of full-sequence pooling throughout the paper: it is not a conservative approximation that dilutes the audio signal, but rather the appropriate strategy given how transformers distribute information.
@@
Table 9: Probe accuracy with all-token vs. audio-only pooling at LLM layers (Ultravox). All deltas are within $\pm$2.1%, most under 1%. Audio-only pooling does not improve probe accuracy. The LLM’s self-attention layers mix information across all sequence positions: by layer 16, audio-derived information has been redistributed to text prompt tokens as well. Restricting pooling to audio positions therefore discards information that the LLM has already spread across the full sequence, slightly reducing the available signal. This validates our use of full-sequence pooling throughout the paper.
Appendix C Additional figures
Figure 1: Probe accuracy trajectories for all five models. Rows 1–2: speech models (Ultravox, Qwen2-Audio) across LibriSpeech, CREMA-D, and ESC-50. Row 3: vision models (LLaVA, Prismatic-DINOv2, Prismatic-SigLIP) on COCO.
Figure 2: Mode alignment profiles for all five models. Blue: alignment score $\tilde{\alpha}(u_{k})$; orange: eigenvalue spectrum $\lambda_{k}$ (log scale). The dominant eigenmodes are modality-specific ($\tilde{\alpha}\approx 0$) for all models with non-text-aligned encoders; for Prismatic-S (SigLIP, text-aligned), even Mode 0 is text-aligned ($\tilde{\alpha}=0.83$).
@@ Appendix C Additional figures
The mode alignment spectrum (Table 11) reveals which directions carry the mismatch. For each model, we decompose the adapter output covariance into eigenmodes and measure their text alignment. The dominant eigenmode is modality-specific ($\tilde{\alpha}\ll 1$) for all non-text-aligned encoders, while for SigLIP even the dominant mode is text-aligned ($\tilde{\alpha}=0.83$).
Model Mode 0 $\tilde{\alpha}$ Mean $\tilde{\alpha}$ (top 10) MS modes MS var.
@@
Table 11: Mode alignment summary. MS modes = modes with $\tilde{\alpha}<0.5$ (modality-specific). MS var. = fraction of top-100 eigenmode variance in MS modes. Mode 0 is modality-specific ($\tilde{\alpha}\ll 1$) for all non-text-aligned encoders; for SigLIP (text-aligned), even Mode 0 is text-aligned ($\tilde{\alpha}=0.83$). Extending to all 4096 modes strengthens the pattern: tail modes are overwhelmingly text-aligned with negligible individual variance, so the MS/TA split is driven by the high-variance modes reported here.
Appendix D Additional limitations

### v2

@@ A.2 GMI-Wasserstein bound
Next we bound the degradation of GMI when the decoder operates on the modal representations $P_{\!M}$ rather than the text representations $P_{\!T}$ it was trained on. The basic idea is as follows: if modal representations are sufficiently close to text representations, then the decoder should achieve near the same performance on them. How “sufficiently close” is determined by two properties of the decoder: how sensitive the decoder is to small perturbations of its inputs, and how large the space of representations is. Theorem 2 relies on three assumptions: the shared marginal (Assumption 1), the Lipschitz condition on the decoder (Assumption 2), and bounded support (Assumption 3).
Assumption 1 (Shared marginal).
The text and modal laws have the same prompts and target tokens: $P_{\!M}(C,Y)=P_{\!T}(C,Y)$.
@@
There exists $L_{\log}>0$ such that for all $(c,y)$ and all $z,z^{\prime}$ in the typical-state region $\mathcal{Z}$: $|\ell(c,z,y)-\ell(c,z^{\prime},y)|\leq L_{\log}\|z-z^{\prime}\|$.
That is, perturbing the representation by any amount changes the log-probability assigned by the decoder by at most $L_{\log}$ times that amount. $L_{\log}$ encodes the sensitivity of the decoder to changes to the input. Neural networks with smooth activations meet this condition locally; we measure $L_{\log}$ empirically using gradient norms (§5.6).
@@
This means the representations that are actually used in practice are confined to a bounded region, rather than being distributed throughout $\mathbb{R}^{d}$. This is precisely what LayerNorm accomplishes: all hidden states are projected onto a hypersphere of radius $\sqrt{d}$.
@@
$|\mathrm{GMI}_{P_{\!M}}(q)-\mathrm{GMI}_{P_{\!T}}(q)|\leq\big(1+e^{L_{\log}\cdot D}\big)\cdot L_{\log}\cdot\mathbb{E}_{(C,Y)}\!\left[W_{\!1}\!\big(P_{\!M}(Z|C,Y),\,P_{\!T}(Z|C,Y)\big)\right].$ (4)
Intuition.
@@
The difference between how much information the decoder can get from text representations and how much it can get from modal representations is that distance between distributions, multiplied by the strength of the decoder. But we have $W_{\!1}$ of the distributions, not the distance between the means; getting the mean of the modal distribution to the mean of the text distribution is not enough. If the modal distribution has a different shape, or spread, or tails, then the exponential penalty factor $(1+e^{L_{\log}D})$ gets them. This is why the text-aligned encoders help to bound (they project on text-like directions, so they decrease both the distance and the shape mismatch), but mean centering would not.
The prefactor $(1+e^{L_{\log}D})$ comes from the “competition term” in the proof (below). In the ambient space this prefactor is trivially large ($L_{\log}\cdot D$ ranges from 10 to 150) and the support-restricted bound (§A.3) refines it in terms of the effective support diameter.
@@ A.2.1 Proof of Theorem 2
Proof overview. There are two terms in the GMI: (A) the expected score of the true representation, and (B) the expected score of a random representation. We must prove that the quantity in both (A) and (B) change very little when we replace the text representations with the modal representations. For (A), this is a corollary of the Lipschitz assumption. For (B), we have a potential complication of an exponential of the scores. However, the bounded-diameter assumption ensures that representations are all contained in a bounded region, which prevents large changes due to the exponential.
Setup. We denote $\ell(z)\equiv\ell(c,z,y)$ for fixed $(c,y)$. The GMI functional separates as $\Gamma(P)=A(P)-B(P)$ where:
@@
We derive upper bounds for $|A(P_{\!M})-A(P_{\!T})|$ and $|B(P_{\!M})-B(P_{\!T})|$ separately.
Step 1: Direct term (how much does the correct representation’s score change?). We would like to bound how much the expected score changes when we go from $P_{\!T}$ to $P_{\!M}$. The Kantorovich–Rubinstein theorem (Kantorovich and Rubinstein, 1958; Villani, 2009) states that: if a function changes by at most L when you change its input by 1 unit ($L$-Lipschitz), then the difference in its expectation w.r.t. two distributions is at most L times the Wasserstein distance between the two distributions. Since $\ell$ is $L_{\log}$-Lipschitz (Assumption 2):
@@
If modal and text representations are similar (small $W_{\!1}$), and the decoder scores are smooth (small $L_{\log}$), then the expected score changes by a tiny amount.
Step 2: Competition term (how much does the random-alternative score change?). This is harder because $B(P)=\log\mathbb{E}[e^{\ell(Z)}]$ involves an exponential of the score, which means small differences are magnified. We can use a coupling $\pi$: a way of pairing up samples from $P_{\!M}$ and $P_{\!T}$ (think of arranging modal representations alongside the text representations they resemble the most). For each pair $(Z_{M},Z_{T})$, define $\Delta\ell=\ell(Z_{M})-\ell(Z_{T})$ as the difference in score between paired representations. By the Lipschitz assumption, $|\Delta\ell|\leq L_{\log}\cdot\|Z_{M}-Z_{T}\|$, and by the bounded diameter (Assumption 3), $|\Delta\ell|\leq L_{\log}\cdot D$.
@@
The first term, $e^{\ell(Z_{T})}/m(P_{\!T})$, is a probability density under $P_{\!T}$ (it sums to 1). The second term, $e^{\Delta\ell}$ accounts for the change in the score when moving from $Z_{T}$ to $Z_{M}$. Since $|e^{x}-1|\leq|x|\cdot e^{|x|}$ for bounded $|x|\leq L_{\log}D$:
@@
Using $|\log(1+x)|\leq|x|$ for $|x|<1$ (applies when $W_{\!1}$ is small enough that the ratio stays near 1):
@@
The $e^{L_{\log}D}$ term is the cost of the exponential: the competition term is more unstable to distributional shift than the direct term, because the former contains $e^{\ell}$ while the latter contains $\ell$.
Step 3: Combine. The overall GMI change is at most the sum of the direct and competition terms:
@@
The competition term dominates: it is $e^{L_{\log}D}$ times larger than the direct term. This is why the prefactor in the bound is so large and why the support-bounded corollary (which reduces $D$) is of practical value.
Step 4: Average over $(C,Y)$. Everything above was for a fixed prompt $C$ and target $Y$. By Assumption 1 ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), the same prompts and targets appear under both laws, so we can simply average over all $(C,Y)$ pairs to obtain the final bound.
When is the bound informative? The bound is a perturbation result: it is tightest when the mismatch $W_{\!1}$ is moderate relative to the decoder’s sensitivity. In the large-mismatch regime ($L_{\log}\cdot W_{\!1}\gg 1$), the prefactor $(1+e^{L_{\log}D})$ makes the right-hand side exceed the maximum possible GMI, and the bound becomes vacuous. In this regime, the bound correctly predicts that collapse occurs (the GMI ceiling is far below the matched-decoder rate) but cannot quantify how much information survives. Empirically, speech models operate in this regime ($L_{\log}\cdot W_{\!1}=19$$162$), which is why the bound is loose for speech but informative for vision ($L_{\log}\cdot W_{\!1}=13$$54$). The support-restricted corollary below partially addresses this by replacing the ambient diameter $D$ with the smaller effective diameter $D_{\text{eff}}$, reducing the prefactor from $\sim e^{150}$ to $\sim e^{7}$.
@@ A.3 Support-restricted bound
Support-restricted bound. If $P_{\!M}(Z|c,y)$ and $P_{\!T}(Z|c,y)$ are supported on some subset $\mathcal{S}\subseteq\mathcal{Z}$ such that $D_{\mathcal{S}}=\mathrm{diam}(\mathcal{S})\leq D$, then $D$ may be replaced by $D_{\mathcal{S}}$ in Theorem 2.
Proof. If both marginals are supported in $\mathcal{S}$, any coupling $\pi$ will also be supported on $\mathcal{S}\times\mathcal{S}$, so $\|Z_{M}-Z_{T}\|\leq D_{\mathcal{S}}$ for $\pi$-a.e. $(Z_{M},Z_{T})$. Step 2 above then gives that $|\Delta\ell|\leq L_{\log}\cdot D_{\mathcal{S}}$, which is the improved constant factor.
Why this matters. The ambient diameter $D$ makes the prefactor $(1+e^{L_{\log}D})$ extremely large ($L_{\log}\cdot D$ ranges from 10 to 150), while representations in modal and text space frequently lie in a far smaller effective volume. We estimate the effective support diameter $D_{\text{eff}}$ from the participation ratio of the pooled representation covariance (§5.6), yielding $L_{\log}\cdot D_{\text{eff}}\in[3,7]$ and a moderate prefactor ($\sim$20–1100).
A.4 Per-type prediction
The theory predicts the overall trend: text-correlated information types (lexical) should strengthen or hold steady through the LLM, while non-text types (speaker, emotion, acoustic) should weaken. This qualitative prediction holds across all five models.
A stronger prediction would be that the degree of weakening for each information type matches its share of modality-specific variance. This does not hold at the level of individual types (Spearman $\rho=-0.40$, $p=0.60$ across the four speech information types in Ultravox). The likely cause is the content-dependent $L_{\log}$: the Lipschitz constant varies $2.5\times$ across audio types ($L_{\log}=14.58$ for speech vs. $5.92$ for environmental sounds in Ultravox), acting as an additional driver of degradation independent of geometric alignment. A per-type bound would need to incorporate this content-dependence, replacing the scalar $L_{\log}$ with a type-conditional $L_{\log}(\tau)$.
A.5 Probe-decoder asymmetry
@@
Proof. Identical to Step 1 of the Theorem 2 proof, replacing $\ell$ with $\log h$ and $L_{\log}$ with $L_{h}$. Also there is no competition term for probes (no random-coding argument), so the exponential prefactor does not appear.
The key asymmetry. A linear probe is a simple model: one matrix multiplication followed by a softmax. Its sensitivity to distributional shift ($L_{h}$) is therefore simply the norm of the weights of that matrix. The LLM decoder is a deep transformer with billions of parameters, and its sensitivity ($L_{\log}$) reflects that depth. Empirically, $L_{\log}/L_{h}\approx 30\times$5.6).
Now consider the same distributional shift with modal representations replacing text representations. The probe’s bound says degradation $\leq L_{h}\cdot W_{\!1}$ (small). The decoder’s bound says degradation $\leq(1+e^{L_{\log}D})\cdot L_{\log}\cdot W_{\!1}$ (large). Same shift, different impact. This is the formal basis for the “present but inaccessible” phenomenon: a probe can still read the information because it is a simple, low-sensitivity instrument, but the decoder’s complex scoring rule is far more sensitive to the distributional mismatch and cannot use it.
@@ B.2 Lipschitz estimation
$L_{\log}$ estimated via the 2-nearest-neighbor gradient norm method over n=1,000n=1{,}000 samples. For each sample $z_{i}$, we compute $\|\nabla_{z}\log q(y_{i}|z_{i},c_{i})\|$ via autograd. The 95th percentile of the resulting distribution serves as the $L_{\log}$ estimate.
@@ B.3 Extraction prompts
For the representation extraction experiments, we provide each model with a minimal, task-agnostic input to prevent the hidden states from being biased toward a specific downstream task. The representations are always extracted from the intermediate hook locations, not from the output.
@@ B.4 LoRA experiment details
LoRA configuration. Rank r=16r{=}16, $\alpha{=}32$, dropout $0.05$, target modules: q_proj, k_proj, v_proj, o_proj.
@@
We calculate the loss only on the assistant response tokens, while the input/prompt tokens are masked (with label 100-100).
@@
Baseline evaluation. Base Ultravox (no LoRA) achieves $17.3\%$ task accuracy on the held-out split (near chance for 6 classes), defaulting to “neutral” for most inputs.
@@ B.5 Full probe accuracy tables
Table 5 reports the full probe accuracies (mean $\pm$ std over 5 seeds) at all four hook points for every model–dataset–information-type combination. The main finding is that probe accuracy for non-text information types remains well above chance even at the LLM’s final layer, confirming that the information survives the decoder’s processing. It is the decoder’s use of this information, not its presence, that fails.
@@ B.6 Information retention overview
Table 6 summarizes the information retention pattern across all five models. Retention is computed as the ratio of probe accuracy at the LLM’s final layer to probe accuracy at the adapter output, normalized by chance. Values above 100% indicate the LLM amplifies the signal; values below 100% indicate degradation. The pattern to notice is that non-text aligned encoders exhibit an asymmetric pattern (lexical content is amplified while speaker identity degrades), while text aligned encoders preserve or improve all the information types, simply because the encoder has already filtered out non-text information.
@@ B.9 Audio-token-only pooling
Table 9 shows the result: audio-only pooling produces nearly identical probe accuracies across all datasets and information types, with all deltas within $\pm$2.1% and most under 1%. Restricting to audio tokens does not help, and in several cases slightly hurts. This is consistent with the LLM mixing information across sequence positions during processing, so that the audio-derived signal is available at non-audio positions by layer 16.
@@
Table 9: Probe accuracy with all-token vs. audio-only pooling at LLM layers (Ultravox). All deltas are within $\pm$2.1%, most under 1%. Audio-only pooling does not improve probe accuracy.
Figure 1: Linear probe accuracy at four extraction points (encoder, adapter, LLM layer 16, LLM final) for all five models. For speech models (rows 1–2), lexical accuracy increases through the LLM while speaker identity degrades, illustrating the information accessibility gap. For vision models (row 3), LLaVA and Prismatic-S (text-aligned encoders) show improvement across all information types, while Prismatic-D (non-aligned) shows flat or declining non-textual accuracy. Datasets: LibriSpeech, CREMA-D, ESC-50 (speech); COCO (vision).
Figure 2: Mode alignment profiles for all five models. Each panel shows the top modes (eigenvectors of the adapter-output covariance $\Sigma_{M}$, sorted by decreasing eigenvalue; 100 for speech models, 200 for vision). Blue (left axis): alignment score $\rho_{k}$, the ratio of text variance to modal variance along each mode (§5.3). Orange (right axis): eigenvalue $\lambda_{k}$ (log scale), the amount of adapter variance carried by each mode. Annotations show the percentage of total variance in Mode 0 and the count of modality-specific modes ($\rho_{k}<0.5$). For non-aligned encoders, the highest-variance modes have $\rho_{k}\approx 0$: the adapter’s dominant variation is along directions the decoder has never seen. For Prismatic-S (SigLIP, text-aligned), even Mode 0 is text-aligned ($\rho_{0}=0.83$).
@@
The mode alignment spectrum (Table 11) reveals which directions carry the mismatch. For each model, we decompose the adapter output covariance into eigenmodes and measure their text alignment. The dominant eigenmode is modality-specific ($\rho_{k}\ll 1$) for all non-text-aligned encoders, while for SigLIP even the dominant mode is text-aligned ($\rho_{k}=0.83$).
Model Mode 0 $\rho_{k}$ Mean $\rho_{k}$ (top 10) MS modes MS var.
@@
Table 11: Mode alignment summary. MS modes = modes with $\rho_{k}<0.5$ (modality-specific). MS var. = fraction of top-100 eigenmode variance in MS modes. Mode 0 is modality-specific ($\rho_{k}\ll 1$) for all non-text-aligned encoders; for SigLIP (text-aligned), even Mode 0 is text-aligned ($\rho_{k}=0.83$). Extending to all 4096 modes strengthens the pattern: tail modes are overwhelmingly text-aligned with negligible individual variance, so the MS/TA split is driven by the high-variance modes reported here.
Appendix C Additional limitations